OpenStudy (anonymous):

What is the solution to the equation 7(a – 10) = 13 – 2(2a + 3)?

OpenStudy (yanasidlinskiy):

Do you know how to use Distributive Property?

OpenStudy (anonymous):

7a -70 = 13 - 4a - 6 7a - 70 = 7 - 4a 11a = 77 a = 7

OpenStudy (anonymous):

7a-70=13-4a-6 11a=77 a=7

OpenStudy (anonymous):

no my quiz is on solving equations and ive been trying to plug in the answers but i still get them wrong

OpenStudy (yanasidlinskiy):

What are your answer choices??

OpenStudy (anonymous):

The answer is satisfying. I checked

OpenStudy (anonymous):

yes it was right. if i could learn how to solve these simple equations i would have it.

OpenStudy (anonymous):

=7

OpenStudy (yanasidlinskiy):